# Exploring a New Design Paradigm for Omnidirectional MAVs for Minimal Actuation and Internal Force Elimination: Theoretical Framework and Control

### v1

@@ Abstract
This paper presents a novel concept for achieving omnidirectionality in a multirotor aerial vehicle (MAV) that uses only 6 inputs and ensures no internal forces at the equilibria. The concept integrates a single actively-tilting propeller along with 3 pendulum-like links, each carrying a propeller, connected by passive universal joints to the main body. We show that this design ensures omnidirectionality while minimizing the internal forces and without resorting to overactuation (i.e., more than 6 inputs). A detailed dynamic model of the multi-link MAV is first developed. Afterwards, the analysis identifies the equilibrium configurations and illustrates that a forced equilibrium exists for every pose of the MAV’s main platform. In order to render this equilibrium asymptotically stable for the closed-loop system, a geometric nonlinear controller is constructed using dynamic feedback linearization and backstepping techniques with the main platform configuration error being the left-trivialized error on $SE(3)$. The stability of the closed-loop system is then investigated by employing standard Lyapunov arguments on the zero dynamics. We conclude by providing numerical simulations validating the proposed approach. They demonstrate the MAV capability to perform decoupled attitude and translational motions under non-zero initial conditions, parametric uncertainty, and actuators noise.
@@ I-B Contributions
The main contribution of this work is to propose the first MAV design, to the best of the authors’ knowledge, having the following properties in SE(3).
@@
    Stabilizability property. We prove in Theorem 1 that it is possible to create a geometric state feedback law that renders the closed-loop system of the MAV asymptotically stable at any base pose in $SE(3)$, as long as configuration and thrust trajectories of the system remain in a certain neighborhood of that base pose. We observe that the dynamics of the passive propeller joints coincide with the zero dynamics, and we prove its asymptotical stability using the Lyapunov theory. The feedback control law is given in Proposition 2.
@@ I-D Outline
The rest of the paper is organized as follows. Sec. II describes the dynamic model of the proposed novel MAV conceptual idea, whereas Sec. III identifies the equilibria sets of that model being all the poses in SE(3) and all having the internal force property. Afterwards, we investigate and demonstrate that the system can be made asymptotically stable at all these equilibria by a state feedback in Sec. IV. Numerical simulations where the MAV performs decoupled motions are the subject of Sec. V. We conclude by giving some remarks on the future work in Sec. VI. For ease of exposition, all the formal proofs are collected in a separate section: Sec. VII.
@@ II Proposed Design and Dynamical Model
In this section, we describe the proposed design and derive its dynamical model in state-space representation.
@@ II-A Mechanical Design Idea
The vehicle concept, depicted in Fig. 1, is composed of a main rigid body called the base and denoted with $B_{1}$, and four smaller bodies, referred to as links and denoted with $B_{2},B_{3},B_{4}$ and $B_{5}$. Each link is connected to the base by means of two revolute joints, thoroughly described in the following, and hosts a rigidly attached velocity controlled brushless (BL) motor driving a propeller. We refer as the actuation unit to each sub-system composed by the propeller, the BL motor, the link, and the two revolute joints connecting the link with the base, for a total of four actuation units. The rotation axes of the two revolute joints of each actuation unit are assumed to intersect in one point, called the center of rotation of the corresponding link, which also coincides with the center of the propeller.
The CoM’s of $B_{2}$, $B_{3}$, and $B_{4}$ are located substantially below their centers of rotation. This can be achieved for example by lowering enough the position of the BL motors of these actuation units, which typically are the heaviest parts. This design makes the three links resemble three spatial pendula with the corresponding joints acting as their pivots. The CoM of $B_{5}$ is instead assumed to be coincident with its center of rotation, for a reason that will be clear in the following.
The actuation units are mounted on the base in such a way that their centers of rotation (i.e., the propeller thrust application points) are all placed on the same plane. This is referred to as a coplanar design.
The two revolute joints connecting $B_{5}$ with $B_{1}$ are assumed to be actuated by two servomotors, representing together a universal joint denoted with $J_{15}$, whose configuration is denoted with $q_{5}\in\mathbb{T}^{2}$. The six revolute joints connecting $B_{1}$ with $B_{2}$, $B_{3}$, and $B_{4}$, are passive, representing three universal joints denoted with $J_{12}$, $J_{13}$, and $J_{14}$, respectively, whose configurations are denoted with $(q_{2},q_{3},q_{4})\in\mathbb{T}^{6}$.
For what has been said so far, the generic universal joint $J_{1j}$, with $j=2,\ldots,5$, is then constituted by two consecutive 1-DoF mutually-orthogonal revolute joints whose configurations are represented by $q_{j1}\in\mathbb{S}^{1}$ and $q_{j2}\in\mathbb{S}^{1}$. We assign body-fixed reference frames to $B_{1},B_{2},\ldots,B_{5}$ with origins coincident with the corresponding CoM’s, and we denote them with $\mathcal{F}_{b}$, $\mathcal{F}_{p2}$, $\mathcal{F}_{p3}$, $\mathcal{F}_{p4}$ and $\mathcal{F}_{p5}$ respectively. The rotational axes corresponding to $q_{j1}$ and $q_{j2}$ are given by the unit vectors, expressed in frame $\mathcal{F}_{b}$, $e_{j1}=(\,\,1\,\,0\,\,0\,\,)^{T}$ and $e_{j2}=R_{x_{b}}(q_{j1})\,(\,\,0\,\,1\,\,0\,\,)^{T}$, respectively, where $R_{k}(*)$ is a pure rotation around axis $k$ by the angle *. Moreover, we assume that each joint composing the passive universal joint $J_{1j}$, with $j=2,3,4$, has viscous friction torques applied about $e_{j1}$ and $e_{j2}$ and symbolized by $\tau_{f_{j1}}$ and $\tau_{f_{j2}}$, respectively.
For the sake of model simplicity, the actuated joint $J_{15}$ is assumed to be driven by two fast servo DC motors taking the joint velocity $u_{v}\in\mathbb{R}^{2}$ as a control input. The motor dynamics is assumed negligible. Furthermore, to improve the formal tractability of the model, we assume that the mass distribution of $B_{5}$ is concentrated at its CoM point which is coincident with its center of rotation. Under such convenient assumptions, the angular moment of inertia of link $B_{5}$ referred to $\mathcal{F}_{p5}$ is zero. This causes the dynamics of the link $B_{5}$ projected onto the minimal coordinates $q_{5}$ to be exactly zero due to the fact that both the time rate of change of the body momentum along with the gravitational wrench vanish when mapped through the screw encoding the kinematical constraints between $B_{1}$ and $B_{5}$. Therefore, the dynamics of $q_{5}$ is solely governed by the motor rotational speed, i.e., $\dot{q}_{5}=u_{v}$
@@
We differentiate between two types of Passive Joints Designs (PJD), denoted in the following as PJD1 and PJD2, respectively. The design PJD1 has no additional feature with respect to the one described so far, whereas the design PJD2 includes an additional Pulley-Belt-Spring module as illustrated in Fig. 2. This schematic shows a pulley $(p_{j}1)$ transferring the rotation around the axis $\bar{y}_{p_{j}}$, attached to the link ${B}_{j}$, with $j=2,3,4$, by an amount of $q_{j2}$, to equivalent rotation around the axis $\bar{x}_{b}$, attached to the intermediate link $\bar{B}_{j}$, through the belt $(b)$ and pulley $(p_{j}2)$. The torsion spring $(s)$, wound along the axis $\bar{x}_{b}$, is connected to the pulley $(p_{j}2)$ such that it compresses with the rotation of this pulley. The spring is fixed at the other end to a point in $\bar{B}_{j}$.
The spring thus applies a moment to the link $\bar{B}_{j}$ around $\bar{x}_{b}$ that is a function of $q_{j2}$. The entire structure of PJD2 moves with $\bar{B}_{j}$, hence the rotation of joint $q_{j1}$ between ${B}_{1}$ and $\bar{B}_{j}$ has no effect on the operation of the Pulley-Belt-Spring module whatsoever. In other words, the mechanism is unidirectional, meaning that rotation of ${B}_{j}$ around $e_{j2}$ causes a moment applied to $\bar{B}_{j}$ around $e_{j1}$ while the reverse is not true due to the module being fixed in place w.r.t $\bar{B}_{j}$ and housed in it. The rationale behind introducing PJD2 and how the spring is tuned shall become clear in Sec III-A. The analysis presented herein applies to both cases, unless stated otherwise explicitly.
@@ II-B Equations of Motion
We denote with $\mathcal{Q}=(x_{1},q_{5})\in SE(3)\times\mathbb{T}^{8}$ the configuration of the whole system, where $x_{1}\in SE(3)\times\mathbb{T}^{6}$ is defined as
$x_{1}=\left(\begin{array}[]{c}x_{11}\\ x_{12}\end{array}\right):=\left(\begin{array}[]{c}C_{1}^{0}\\ q_{t}\\ \end{array}\right)$ (1)
where $x_{11}:=C_{1}^{0}\in SE(3)$ is the configuration of the base $B_{1}$ relative to the inertial world frame $\mathcal{F}_{w}$ whereas $x_{12}:=q_{t}=(q_{2},q_{3},q_{4})\in\mathbb{T}^{6}$. The velocity of $x_{1}(t)$ is denoted by $x_{2}\in se(3)\times\mathbb{R}^{6}$ which is given as
$x_{2}:=\left(\begin{array}[]{c}x_{21}\\ x_{22}\end{array}\right)=\left(\begin{array}[]{c}V_{1}^{0}\\ \dot{q}_{t}\end{array}\right)$ (2)
where $x_{21}:=V_{1}^{0}=\dot{C}^{0}_{1}(C^{0}_{1})^{-1}\in se(3)$ represents the right differential on $SE(3)$, or equivalently, the right invariant (spatial) twist of the base $B_{1}$ relative to, and expressed in, the inertial world frame $\mathcal{F}_{w}$, while $x_{22}:=\dot{q}_{t}=(\dot{q}_{2},\dot{q}_{3},\dot{q}_{4})\in\mathbb{R}^{6}$ are the joint speeds of the passive joints.
The motion is then described by the dynamical equations
$\begin{split}&M(x_{1})\dot{x}_{2}+h(x_{1},{x}_{2})+\bar{g}(x_{1})-\bar{\tau}_{f}(x_{22})=W^{0}_{app}(x_{1})\\ &\dot{q}_{5}=u_{v}\end{split}$ (3)
where the meaning of all the symbols not yet defined and appearing in (3) are provided in Table I with $j=2,3,4$ and $i=1,2$.
Symbol Definition Formula
$q_{j}$ Joint $J_{1j}$’s configuration $q_{j}=\operatorname{col}(q_{ji})\in\mathbb{T}^{2}$
$q_{t}$ Passive joints configuration $q_{t}=\operatorname{col}(q_{j})\in\mathbb{T}^{6}$
$S_{j}$ Joint $J_{1j}$’s spatial instantaneous screw $S_{j}=(\,S_{j1}\;S_{j2}\,),\ \text{with }S_{ji}\in\mathfrak{se}(3)$
$S_{t}$ Passive joints spatial instantaneous screw list $S_{t}=\operatorname{blockdiag}(S_{j})$
$M$ Generalized mass matrix $M=\begin{pmatrix}M_{bb}&M_{bt}\\ M_{bt}^{T}&M_{tt}\end{pmatrix}$
$M_{bb}$ Generalized base $B_{1}$ mass matrix $M_{bb}=M^{0}_{B_{1}}+\sum^{5}_{j=2}M^{0}_{B_{j}}$, with $M^{0}_{B_{j}}$ being $B_{j}$’s inertia tensor referred to the world frame $\mathcal{F}_{w}$
$M_{bt}$ Generalized coupling mass matrix $M_{bt}=\sum^{4}_{j=2}M^{0}_{B_{j}}S_{j}$
$M_{tt}$ Generalized joint mass matrix $M_{tt}=\operatorname{blockdiag}(S_{j}^{T}M^{0}_{B_{j}}S_{j})$
$\dot{M}^{0}_{j}$ Time derivative of $B_{j}$’s inertia tensor referred to the world frame $\mathcal{F}_{w}$ $\dot{M}^{0}_{j}=-M^{0}_{j}\operatorname{ad}_{V^{0}_{j}}-\operatorname{ad}^{T}_{V^{0}_{j}}M^{0}_{j}$, with $\operatorname{ad}_{V^{0}_{j}}$ denoting the adjoint operator on $V^{0}_{j}$ (the spatial twist of body $B_{j}$), $V^{0}_{j}=V^{0}_{1}+S_{j}\dot{q}_{j}\ \forall j\in\{2,3,4\}$. Also, $\dot{M}^{0}_{5}=-M^{0}_{5}\operatorname{ad}_{V^{0}_{1}}-\operatorname{ad}^{T}_{V^{0}_{1}}M^{0}_{5}$.
$\dot{S}_{j}$ Time derivative of screw $S_{j}$ $\dot{S}_{j}=\operatorname{ad}_{V^{0}_{j}}S_{j}$
$h$ Coriolis & centrifugal terms $h=\big(h_{b}\;\;\operatorname{col}(h_{j})\big)^{T}$
$h_{b}$ Base $B_{1}$ Coriolis & centrifugal terms $h_{b}=\dot{M}_{bb}V^{0}_{1}+\sum^{4}_{j=2}\big(M^{0}_{B_{j}}\dot{S}_{j}+\dot{M}^{0}_{B_{j}}S_{j}\big)\dot{q}_{j}$
$h_{j}$ $B_{j}$’s Coriolis & centrifugal terms projected onto joint $J_{1j}$ $h_{j}=S_{j}^{T}\big((M^{0}_{B_{j}}\dot{S}_{j}+\dot{M}^{0}_{B_{j}}S_{j})\dot{q}_{j}+\dot{M}^{0}_{B_{j}}V^{0}_{1}\big)$
$W_{grav,B_{j}}^{0}$ Spatial gravitational wrench applied at $B_{j}$’s CoM $W_{grav,B_{j}}^{0}=M_{B_{j}}^{0}G^{0}$, with $G^{0}=\operatorname{col}(0_{5},-g)$. Here $g$ is the gravity acceleration.
$\bar{g}$ Gravitational terms $\bar{g}=\big(g_{b}\;\;g_{t}\big)^{T}$
$g_{b}$ Base $B_{1}$ gravitational force $g_{b}=-\sum^{5}_{j=1}W_{grav,B_{j}}^{0}$
$g_{t}$ Joint gravitational torque $g_{t}=-S_{t}^{T}\operatorname{col}(W_{grav,B_{j}}^{0})$
$W^{0}_{app,b}$ Applied external spatial wrench at base $B_{1}$’s CoM $W^{0}_{app,b}=\sum^{5}_{j=2}W_{app,B_{j}}^{0}\in\mathfrak{se}(3)^{*}$, with $W_{app,B_{j}}^{0}\in\mathfrak{se}(3)^{*}$ being the applied spatial wrench at $B_{j}$’s CoM
$W^{0}_{app,t}$ Applied external spatial wrench at pendulum links $B_{2},B_{3},B_{4}$ projected onto the respective joints through $S_{t}$ $W^{0}_{app,t}=S_{t}^{T}\operatorname{col}(W_{app,B_{j}}^{0})$
$W^{0}_{app}$ Total applied external wrenches $W^{0}_{app}=\big(W^{0}_{app,b}\;\;W^{0}_{app,t}\big)^{T}$
$\tau_{fj}$ Joint $J_{1j}$’s applied viscous friction torque $\tau_{fj}=\operatorname{col}(\tau_{f_{ji}})\in\mathbb{R}^{2}$, and $\bar{\tau}_{f}=\big(0_{6\times 1}\;\;\operatorname{col}(\tau_{f_{j}})\big)^{T}$
TABLE I: Variable definitions and formulas
Denote the full state vector $x(t)$ as $x=(x_{1},x_{2},{q}_{5})^{T}\in\mathbb{X}\equiv SE(3)\times\mathbb{T}^{8}\times se(3)\times\mathbb{R}^{6}$. Then (3) can be put in the state space form $\Sigma_{1}$ as follows
@@
where $[.]$ denotes the matrix form of $se(3)$, and $D_{b}$, $D_{t}$ and the input vector $u$ are defined as
$\begin{split}D_{b}=&\left(\begin{array}[]{cccc}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}&\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}\end{array}\right)_{6\times 4}\\ D_{t}=&\left(\begin{array}[]{ccc}S_{2}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&0&0\\ 0&S_{3}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&0\\ 0&0&S_{4}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}\end{array}\right)_{6\times 3}\\ u=&\left(\begin{array}[]{cccccc}f_{2}&f_{3}&f_{4}&f_{5}&u_{v_{1}}&u_{v_{2}}\end{array}\right)_{6\times 1}^{T}\in\mathbb{R}^{6}\end{split}$
@@
The proposed design satisfies the input dimension property. In fact the dimension of the manifold of the configuration of the base frame $C_{1}^{0}$ is six and the dimension of the input is $\operatorname{dim}(u)=6$ (four propeller speeds and two servomotor speeds). In other words, the system is square with respect to the output being $C_{1}^{0}$, the pose of $B_{1}$.

### v2

@@ Abstract
This paper presents a novel concept for achieving omnidirectionality in a multirotor aerial vehicle (MAV) that uses only 6 inputs and ensures no internal forces at the equilibria. The concept integrates a single actively-tilting propeller along with 3 pendulum-like links, each carrying a propeller, connected by passive universal joints to the main body. We show that this design ensures omnidirectionality while minimizing the internal forces and without resorting to overactuation (i.e., more than 6 inputs). A detailed dynamic model of the multi-link MAV is first developed. Afterwards, the analysis identifies the equilibrium configurations and illustrates that a forced equilibrium exists for every pose of the MAV’s main platform. In order to render this equilibrium asymptotically stable for the closed-loop system, a coordinate-invariant nonlinear controller is constructed using dynamic feedback linearization and backstepping techniques with the main platform configuration error being the left-trivialized error on $SE(3)$. The stability of the closed-loop system is then investigated by employing standard Lyapunov arguments on the zero dynamics. We conclude by providing numerical Gazebo simulations, whose videos are available at https://youtu.be/SWIt0CSwNUo, validating our approach. They demonstrate the MAV capability to perform decoupled attitude and translational motions under parametric uncertainty and actuators noise.
Nomenclature
$SE(3)$
Special Euclidean group of rigid-body transformations.
se(3){se}(3)
Lie algebra of $SE(3)$.
$\mathbb{T}^{n}$
$n$-torus manifold (a compact Lie group).
$C^{0}_{j}\in SE(3)$
Configuration of $B_{j}$’s frame $\mathcal{F}_{p_{j}}$ in the world frame $\mathcal{F}_{w}$
$V^{0}_{1}\in\mathbb{R}^{6}$
Right-invariant base twist expressed in, and relative to, $\mathcal{F}_{w}$
[]:[*]: $\mathbb{R}^{6}\to se(3)$ An isomorphism with an inverse $\hat{*}$.
$(*)_{\times}$ A skew-symmetric matrix of vector $\in\mathbb{R}^{3}$.
$J_{1j}$
Universal joint connecting the base to link $B_{j}$.
$S_{j}\in\mathbb{R}^{6\times 2}$
$J_{1j}$’s spatial instantaneous screw axis, $S_{j}=(\,S_{j1}\;S_{j2}\,)$.
$S_{p}=\operatorname{blkdiag}(S_{j})$
Block-diagonal matrix of screws $S_{j}$.
$\operatorname{Ad}_{C}:$
$SE(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $SE(3)$.
$\operatorname{ad}_{V}:$
$se(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $se(3)$.
$R_{k}(*)$
Pure rotation about axis $k$ by a angle *.
$e_{i}$
zero vector with one in index $i$.
$M$
Generalized mass (inertia) matrix.
$M_{bb},\ M_{bp},\ M_{pp}$
Base, coupling, and joint sub-matrices of $M$.
$h$
Centrifugal and Coriolis terms.
$g_{b},\,g_{p}$
Gravitational terms of base and joints, $g=(\,g_{b}\,\,g_{p}\,)^{T}$.
$W^{0}_{\mathrm{app,b}}$
Applied wrench on base’s CoM expressed in $\mathcal{F}_{w}$.
$W^{0}_{\mathrm{app,p}}$
Applied wrench on links $B_{2}$, $B_{3}$, $B_{4}$’s CoM expressed in $\mathcal{F}_{w}$ and projected onto the respective joints through the corresponding screw in $S_{p}$.
$\tau_{dr_{i}}$
Drag torque about propeller $i$’s axis.
$\tau_{f_{ji}}$
Viscous friction at joint $q_{ji}$’s axis. $\bar{\tau}_{f}=\big(0_{6\times 1}\;\;\operatorname{col}(\tau_{f_{ji}})\big)^{T}$
$\tau_{s_{j}}$
Torsion spring torque applied at joint $q_{j1}$’s axis.
$\tau_{dj}$
Propeller $j$’s drag torque factorized on $q_{j1}$’s axis at equilibrium.
$m_{b},m_{j}$
Mass of base and link $B_{j}$.
$c$
Length from joints axes intersection point to $B_{j}$’s CoM.
$k_{i}$
Drag-to-thrust ratio of propeller $i$.
$\mathrm{PJD}_{1},\ \mathrm{PJD}_{2}$
Passive-joint designs type 1 and type 2.
$P_{se(3)}(A)$
projects a matrix $A\in\mathbb{R}^{4\times 4}$ on $se(3)$.
$_{e}$
Variable * evaluated at equilibrium.
$^{[i]}$
Value of variable * pertaining to joint design PJDi.
@@ I-B Contributions
The main contribution of this work is to propose the first MAV design, to the best of the authors’ knowledge, having the following four properties in SE(3):
@@
    Stabilizability property. We prove in Theorem 1 that it is possible to create a state feedback law that renders the closed-loop system of the MAV asymptotically stable at any base pose in $SE(3)$, as long as configuration and thrust trajectories of the system remain in a certain neighborhood of that base pose. We observe that the dynamics of the passive propeller joints coincide with the zero dynamics, and we prove its asymptotical stability using the Lyapunov theory. The feedback control law is given in Proposition 2.
@@
These properties capture major requirements for enabling energy-efficient omnidirectional flight. In particular, the ability to hover at any attitude in $SO(3)$ is the defining feature of omnidirectional MAVs [14, 15], hence expanding their reachable workspace and manipulation capabilities [12]. Ensuring stabilizability of these equilibria is essential for any practical deployment. Moreover, achieving full 6-DOF control with only 6 actuators alongside eliminating internal forces at hovering directly contributes to significantly improved energy efficiency, simplified input allocation, and extended flight endurance [21, 13, 22]. Together, these properties illustrate the impact of the proposed MAV concept within the field of aerial robotics.
@@ I-D Outline
The rest of the paper is organized as follows. Sec. II describes the dynamic model of the proposed novel MAV conceptual idea, whereas Sec. III identifies the equilibria sets of that model being all the poses in SE(3) and all having the internal force property. Afterwards, we investigate and demonstrate that the system can be made asymptotically stable at all these equilibria by a state feedback in Sec. IV. Numerical simulations where the MAV performs decoupled motions are the subject of Sec. V. We conclude by giving some remarks on the future work in Sec. VI. For ease of exposition, all the formal proofs, derivations, and explicit calculation formulas are collected in appendices of [29].
@@ II Proposed Design and Dynamical Model
In this section, we describe the proposed design and present its dynamical model in state-space form.
@@ II-A Mechanical Design Idea
The vehicle concept, shown in Fig. 1, consists of a main rigid body, referred to as the base and denoted by $B_{1}$, and four smaller bodies, referred to as links and denoted by $B_{2},B_{3},B_{4}$, and $B_{5}$. Each link carries a rigidly-attached motor driving a propeller and is connected to the base by two revolute joints. The propeller, the motor, the link, and the two revolute joints together form an actuation unit. For each unit, the joints’ mutually-orthogonal rotation axes intersect at a point, called the center of rotation of the link, coinciding with the propeller center. The actuation units are mounted on $B_{1}$ such that all centers of rotation lie on the same plane, resulting in a coplanar design. The centers of mass (CoMs) of $B_{2},B_{3}$, and $B_{4}$ are located below their respective centers of rotation, achieved by placing the motors sufficiently low. This makes these links behave as spatial pendula, with the joints acting as pivots. In contrast, $B_{5}$’s CoM coincides with its center of rotation. $B_{1}$ is connected to $B_{5}$ by an actuated universal joint with configuration $q_{a}=\operatorname{col}(q_{5i})\in\mathbb{T}^{2}$, and to $B_{2},B_{3}$, and $B_{4}$ by passive universal joints with configurations $q_{p}=\operatorname{col}(q_{ji})\in\mathbb{T}^{6}$, $j=2,\ldots,5$, $i=1,2$.
For modeling simplicity, the actuated joint is assumed to be driven by fast velocity-controlled servo DC motors, neglecting motor dynamics. Moreover, $B_{5}$’s mass is assumed to be concentrated at its CoM, zeroing its angular inertia moment. Under these assumptions, the projected dynamics of $B_{5}$} onto its $q_{a}$’s screw axis vanish, rendering $\dot{q}_{a}$ governed solely by the motor input.
Two Passive Joint Designs (PJD) are considered. In PJD1, the passive joints are implemented as described above. In PJD2, an additional Pulley-Belt-Spring module (Fig. 2) is integrated into each passive joint. This module transfers the rotation about $q_{j2}$’s axis to a rotation about $q_{j1}$’s axis through a belt-pulley transmission. A torsion spring, mounted along $q_{j1}$’s axis, applies a moment proportional to $q_{j2}$, thus the mechanism is unidirectional: rotations about $q_{j2}$ generate a restoring moment about $q_{j1}$, while the reverse is not true. The rationale behind PJD2 is discussed in Sec. III-A. Unless otherwise stated, the analysis herein applies to both designs.
@@ II-B Equations of Motion
We denote with $\mathcal{Q}=(C_{1}^{0},{\color[rgb]{0,0,0}{q_{p}}},{\color[rgb]{0,0,0}{q_{a}}})\in SE(3)\times\mathbb{T}^{8}$ the system’s configuration. The velocity of $\color[rgb]{0,0,0}{(C_{1}^{0},q_{p})}$} is ${\color[rgb]{0,0,0}{(V_{1}^{0},\dot{q}_{p})}}\in\mathbb{R}^{12}$ where $V_{1}^{0}]=\dot{C}^{0}_{1}(C^{0}_{1})^{-1}\in se(3)$. Denote the full state vector $x(t)$ as $x=({\color[rgb]{0,0,0}{C_{1}^{0},q_{p}}},{\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}},{\color[rgb]{0,0,0}{q_{a}}})^{T}\in\mathbb{X}\equiv SE(3)\times\mathbb{T}^{8}\times\mathbb{R}^{12}$. Dropping dependencies for clarity, the system dynamics is thus described by these state space equations
@@
where explicit formulas of $M,h$ and $g$ are given in Sec. -A [29], and $D_{b}$, $\color[rgb]{0,0,0}{D_{p}}$} and the input vector $u$ are defined as
$\begin{split}D_{b}=&\left(\begin{array}[]{cccc}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}&\operatorname{Ad}^{T}_{(C^{0}_{f_{5}})^{-1}}\bar{f}_{5}\end{array}\right)_{6\times 4}\\ {\color[rgb]{0,0,0}{D_{p}}}=&\left(\begin{array}[]{ccc}S_{2}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{2}})^{-1}}\bar{f}_{2}&0&0\\ 0&S_{3}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{3}})^{-1}}\bar{f}_{3}&0\\ 0&0&S_{4}^{T}\operatorname{Ad}^{T}_{(C^{0}_{f_{4}})^{-1}}\bar{f}_{4}\end{array}\right)_{6\times 3}\\ u=&\left(\begin{array}[]{cccccc}f_{2}&f_{3}&f_{4}&f_{5}&u_{v_{1}}&u_{v_{2}}\end{array}\right)_{6\times 1}^{T}\in\mathbb{R}^{6}\end{split}$
@@
The proposed design satisfies the input dimension property. In fact the dimension of the manifold of the configuration of the base frame $C_{1}^{0}$ is six and the dimension of the input is $\operatorname{dim}(u)=6$ (four propeller thrusts and two servomotor speeds).